A capacitor is a square with an angle a = 3cm with a distance between the ammeters e = 2mm. Calculate the capacitance of the capacitor if the electric field is air.


1
Expert's answer
2021-06-07T09:35:26-0400

Gives

Square side (a)=3cm

Distance between palate (d)=2mm

Area(A)="(3\times10^{-2})^2=9\times10^{-4}m^2"


"C=\frac{A\epsilon_0}{d}"


"C=\frac{9\times10^{-4}\times8.85\times10^{-12}}{2\times10^{-3}}"

"C=3.98\times10^{-12}F"

Electric field

"E=\frac{Q}{A\epsilon_0}"

"E=\frac{Q}{9\times10^{-4}\times8.85\times10^{-12}}"

"E=\frac{Q}{7.97\times10^{-15}}" V/m